# SafeFix: Targeted Model Repair via Controlled Image Generation

### v1

@@ 4.3 Filtering via Large Vision–Language Models
We found that synthetic images generated by the Conditional Diffusion Model (CDM) can sometimes fail to accurately reflect the intended attribute modifications. To address this, we employ large vision–language models (LVLMs), Qwen2.5-VL-7B Bai et al. (2025) and LLaVA-v1.5-7B Liu et al. (2023), to automatically verify that each generated image correctly exhibits the desired attributes and retains the original label.
@@
We retain only those images for which the LVLM answers “yes” to all queries and confirms that the label matches the original ground truth $y(x)$. Thus, after the LVLM filtering, the generated image $x^{\prime}$ satisfies the desired attributes and preserves the original label, i.e., $y(x^{\prime})=y(x)$. These validated samples are then added to $\mathcal{D}_{\mathrm{train}}$ for model retraining.
@@ 4.4 Combining Generated Images with the Original Dataset and Retraining
To repair the rare-case bugs while maintaining the model’s performance on the overall training distribution  Lee et al. (2024), we augment the training set by adding the validated synthetic images $\{x^{\prime}\}$, yielding an updated training set $\mathcal{D}_{\mathrm{train}}^{\prime}=\mathcal{D}_{\mathrm{train}}\cup\{x^{\prime}\}.$ We then retrain the vision model $f_{\theta}$ on $\mathcal{D}_{\mathrm{train}}^{\prime}$ and evaluate its performance on $\mathcal{D}_{\mathrm{val}}$. We report both the overall accuracy improvement and the reduction in failure rates on critical rare-attribute slices $S_{e}$, before and after augmentation.
@@ 5.1 Experimental Setup
Datasets. We evaluate our method on two classification tasks that exhibit attribute-based failure modes.
Lipstick-wearing classification. We use the CelebA dataset Liu et al. (2015) and follow the same train/validation split protocol as Chen et al. (2023), with 80,000 images used for training and 20,000 for validation. The task is to predict whether a person is wearing lipstick, a label known to correlate spuriously with other attributes such as gender and hair color. In the following experiments, we refer to this dataset simply as CelebA.
ImageNet-10 classification. We construct a 10-class subset of ImageNet Deng et al. (2009) containing the following categories: backpack, barber chair, coffee mug, desk, electric guitar, park bench, pitcher, purse, rocking chair, and water bottle. Each class contains 1,300 images, and we use an 8:2 split for training and validation, respectively. This subset is selected to study classification failures related to visual attributes such as texture and color. In the following experiments, we refer to this dataset simply as ImageNet10.
@@
  • DA-CDM  Fang et al. (2024) is a data augmentation method designed specifically for object detection. It utilizes a controllable diffusion model, guiding the generation process by extracting visual priors, such as edge contours, from existing images, which allows for the direct reuse of the original bounding box annotations. Finally, the method employs a category-calibrated CLIP score to filter the generated data, ensuring the high quality and textual relevance of the augmented samples.
  • Mask-ControlNet  Huang et al. (2024) is a text-guided image generation pipeline that uses ControlNet with facial occlusion masks to synthesize diverse face images under specific occlusions as a data augmentation method to improve model robustness.
  • HiBug_Class  Chen et al. (2023). A Class-level method that augments training data with synthetic images. For each class, it uses a diffusion model with a prompt:
@@
  • HiBug_Task Chen et al. (2023). A Task-level variant of HiBug_Class that targets failure-prone attributes. It selects attribute slices with the highest validation error and generates prompts to guide a diffusion model:
@@
    Note: Unless otherwise specified, “HiBug” refers to this optimized “HiBug_Task” variant in the paper. We do not compare with HiBug2 Chen et al. (2025), which is a data selection method that retrieves web images using image search engines, and differs from the synthetic generation strategy.
Implementation Details. All experiments use an NVIDIA A100 GPU. We take the CelebA “wearing lipstick” classification task as an example. We evaluate three backbone architectures for this classification task: ResNet-18 He et al. (2016), ViT-B/16 Dosovitskiy et al. (2020), and CLIP (ViT-B/32) Radford et al. (2021), all of which are initialized with random weights. All vision models are trained using cross-entropy loss. Optimization is performed using Adam Kingma and Ba (2015) with a learning rate of $1\times 10^{-4}$ and weight decay of $1\times 10^{-3}$. All experiments use a batch size of 64 and are trained for 20 epochs.
For rare-case bug discovery and augmentation, we follow the procedure described above, using validation accuracy and slice support to identify candidate attributes. For the main results, we set the rarity threshold $\rho=0.05$ and the accuracy difference threshold $\epsilon=0.01$ based on dataset statistics.
For synthetic augmentation, we use ControlNet Zhang et al. (2023), a CDM based on Stable Diffusion 1.5 Rombach et al. (2022) and conditioned on soft HED boundaries Xie and Tu (2015), with 30 DDIM inference steps. Soft HED boundaries preserve structural details, making this approach suitable for attribute-preserving edits like recoloring and stylizing. To filter generated images, we employ the LVLMs Qwen2.5-VL-7B Bai et al. (2025) and LLaVA-v1.5-7B Liu et al. (2023), ensuring consistent model scaling. Generating 1,000 images with ControlNet takes about one hour, and LVLM filtering takes 10 minutes. Filtering accuracy for most attributes (e.g., hair color, skin tone) exceeds 90%. For emotion-related attributes, Qwen2.5-VL’s accuracy is around 80%, while LLaVA-v1.5’s is over 95%. This discrepancy suggests emotional edits are less reliably generated and harder to verify, indicating mild semantic ambiguity. This results in a 70-80% pass rate for filtered images. Appendix C.2 shows why we choose Qwen2.5-VL in most of our experiments.
@@ 5.2 Main Results
Method ResNet ViT CLIP
Base 90.68 85.07 89.85
DA-CDM 90.71 90.65 90.63 85.39 86.31 85.80 89.91 90.02 89.95
Mask-ControlNet 90.81 91.02 90.91 87.45 87.40 87.88 90.01 90.00 90.07
HiBug_Class 90.87 90.93 90.67 85.65 85.93 85.92 89.86 89.89 89.83
HiBug_Task 90.93 90.87 90.76 87.60 86.57 86.01 90.03 89.98 89.96
Ours (L) 91.32 91.42 91.40 87.49 87.50 87.07 90.71 90.70 90.64
Ours (Q) 91.69 91.38 91.71 87.65 87.07 87.08 90.72 90.73 90.52
Table 1: Validation accuracy (%) on CelebA for varying numbers of added images, models, and methods. Ours (L) and Ours (Q) denote using LLaVA-v1.5-7B and Qwen2.5-VL-7B as the large vision-language model filter, respectively.
@@
On CelebA, we select rare-case bugs defined by attribute–value combinations red hair, brown skin, and sad emotion for ResNet and ViT, and yellow hair, brown skin, and sad emotion for CLIP, based on the most frequent patterns identified among failure slices. SafeFix consistently achieves the highest validation accuracy across all models (ResNet, ViT, and CLIP) and varying levels of synthetic augmentation. For example, with 1,000 added images, our method improves accuracy by +1.01% (ResNet), +2.58% (ViT), and +0.87% (CLIP) relative to their respective baselines. These results show that our attribute-targeted augmentation and filtering pipeline is effective in repairing rare-case failure slices, outperforming both CDM-based and HiBug baselines.
Method ResNet ViT CLIP
Base 71.81 97.46 93.73
DA-CDM 73.05 73.51 73.90 97.42 97.45 97.48 93.85 93.81 93.88
Mask-ControlNet 72.98 72.60 72.11 97.31 97.40 97.35 93.21 93.80 93.75
HiBug_Class 71.98 72.09 72.01 96.95 97.11 97.01 93.75 93.98 93.88
HiBug_Task 73.27 72.73 73.12 97.45 97.56 97.49 93.35 93.92 93.89
Ours (L) 73.76 73.80 74.53 98.00 97.95 98.11 94.19 94.50 94.26
Ours (Q) 74.08 73.12 73.65 98.01 97.96 97.98 94.25 94.52 94.73
Table 2: Validation accuracy (%) on ImageNet10 under varying image counts, models, and methods.
@@
On ImageNet10, similar trends emerge, as shown in Table 2. For ResNet and ViT, we target rare-case bugs involving pink color and fabric texture, while for CLIP we use orange color and fabric texture. Across all models, our proposed method consistently surpasses the baseline methods. Specifically, our method improves ResNet accuracy by +2.27% (at 100 images) compared to the base model. ViT and CLIP also exhibit a steady improvement compared to other methods.
@@
Analysis. Our experiments show that baseline performance is often unstable. For instance, baselines like DA-CDM and Mask-ControlNet show some improvement on object and facial attribute tasks, respectively, due to their use of conditional diffusion models. However, they are not targeted model repair methods, so their overall performance is inferior to both HiBug and SafeFix. This lack of a targeted strategy means their gains are task-specific and not robustly transferable. Similarly, the accuracy of HiBug does not improve substantially. This is likely because it does not precisely target failure-critical attributes, instead applying general augmentations that can introduce noise or modify unintended features. As shown in Figure 2, this issue is compounded by the lack of LVLM filtering, resulting in generated data that is often misaligned with the intended rare-case fixes.
In contrast, our method (SafeFix) consistently shows stable or improving performance. This robustness indicates our attribute-targeted augmentation and LVLM filtering are highly effective at correcting failure-prone subpopulations with meaningful data. However, we also note that in the large-scale setting (e.g., CelebA with 10,000 augmented images), the improvement becomes smaller. One reason may be the relative scale between synthetic and real data: CelebA has 80,000 training images, so adding 10,000 shifts the distribution, but not enough to fully rebalance all rare-case slices. Oversaturation of similar synthetic patterns may also yield diminishing returns. A more detailed analysis is provided in Appendix E.
@@ 5.3 SafeFix Can Effectively Fix Rare-case Bugs
Figure 3: Accuracy comparison of ResNet-18 trained using different augmentation methods. The dashed line represents the average accuracy without additional training data.
To verify that SafeFix’s improvements specifically address targeted rare-case bugs rather than merely enhancing overall performance, we analyze attribute-level accuracy changes on CelebA and ImageNet10, shown in Figure 3. For clarity, the ImageNet10 plot includes only three representative attributes—color, background, and texture—as the dataset contains many attribute dimensions. Specifically, the left plot highlights improvements for CelebA after adding 5,000 synthetic images to the original 80,000 training samples. The right plot demonstrates accuracy gains on ImageNet10, achieved by adding 100 synthetic images to the original training set of 10,400 samples.
Taking ImageNet10 as an example, our targeted synthetic augmentation on pink color and fabric texture significantly improved accuracy for these selected rare-case attributes. Accuracy for the pink color attribute increased from 69.90% to 74.76%, surpassing the overall validation accuracy of 74.08%. Similarly, accuracy for the fabric texture attribute improved from 65.85% to 75.61%, also exceeding the overall accuracy. In contrast, attributes not explicitly targeted by augmentation, such as the rocks background, exhibited minimal or no improvement—its accuracy remained unchanged—highlighting that the gains from SafeFix are concentrated on the intended rare-case bugs rather than uniformly distributed across all attributes.
These substantial attribute-specific improvements confirm that SafeFix effectively repairs identified rare-case failure slices rather than providing a generalized performance boost. Attributes not targeted by augmentation show smaller or negligible accuracy changes, further reinforcing that SafeFix precisely addresses the targeted rare-case failures.
@@ 5.4 SafeFix Directly Addresses Diagnosed Failure Modes, Not Merely Augments Data
To verify that SafeFix ’s performance gains stem from accurately fixing rare-case bugs rather than from generic data augmentation, we compare red-hair and yellow-hair augmentations on CLIP. As discussed in Section 5.2, we select red hair for ResNet and yellow hair for CLIP in the CelebA dataset, based on which attribute is more likely to trigger rare-case bugs.
Figure 4: Accuracy improvements for red-hair vs. yellow-hair augmentation on CLIP.
For CLIP, both red hair and yellow hair are low-frequency attributes that meet the rarity criterion in Eq. (4). However, only the yellow hair slice additionally satisfies the low-accuracy criterion in Eq. (5), making it a true rare-case bug slice for CLIP.
Figure 4 confirms that augmenting with yellow-hair samples—the diagnosed failure mode for CLIP—substantially improves accuracy on its target slice (+5.14%) and provides gains on related attributes. Conversely, augmenting with red hair, which is not a diagnosed bug, reveals a more complex outcome. Most notably, it degrades performance on its own slice by -0.81%. This counter-intuitive result contrasts with the positive gains seen for ResNet (Figure 3(a)), highlighting that different architectures can react to synthetic data in unpredictable ways. While red-hair augmentation does yield some comparable gains on other attributes (e.g., sad emotion and overall), the significant risk of negative side effects on the target class itself underscores the importance of our diagnosis-driven approach for reliable model repair, rather than augmenting indiscriminately.

### v2

@@ 4.3 Filtering via Large Vision–Language Models
We found that synthetic images generated by the Conditional Diffusion Model (CDM) can sometimes fail to accurately reflect the intended attribute modifications. To address this, we employ large vision–language models (LVLMs), Qwen2.5-VL-7B [bai2025qwen25vl] and LLaVA-v1.5-7B [liu2023improvedllava], to automatically verify that each generated image correctly exhibits the desired attributes and retains the original label.
@@
We retain only those images for which the LVLM answers “yes” to all queries and confirms that the label matches the original ground truth $y(x)$. Thus, after the LVLM filtering, the generated image $x^{\prime}$ satisfies the desired attributes and preserves the original label, i.e., $y(x^{\prime})=y(x)$. These validated samples are then added to $\mathcal{D}_{\mathrm{train}}$ for model retraining. We conduct human-audit experiments to verify that the LVLMs can reliably filter out low-quality generated samples.
@@ 4.4 Combining Generated Images with the Original Dataset and Retraining
To repair the rare-case bugs while maintaining the model’s performance on the overall training distribution  [lee2024genmix], we augment the training set by adding the validated synthetic images $\{x^{\prime}\}$, yielding an updated training set $\mathcal{D}_{\mathrm{train}}^{\prime}=\mathcal{D}_{\mathrm{train}}\cup\{x^{\prime}\}.$ We then retrain the vision model $f_{\theta}$ on $\mathcal{D}_{\mathrm{train}}^{\prime}$ and evaluate its performance on $\mathcal{D}_{\mathrm{val}}$. We report both the overall accuracy improvement and the reduction in failure rates(fix rate) on critical rare‐attribute slices $S_{e}$, before and after augmentation.
@@ 5 Results
Method ResNet (base acc: 90.57%) ViT (base acc: 85.02%) CLIP (base acc: 88.32%)
DA-CDM 4.03 3.29 3.08 4.07 10.21 6.81 15.32 16.52 15.58
Mask-ControlNet 5.41 7.32 6.15 11.28 10.48 12.88 16.10 15.92 16.78
HiBug_Class 5.09 5.83 3.40 5.14 7.74 7.21 14.64 14.98 13.96
HiBug_Task 6.79 4.88 3.92 18.69 11.75 7.74 15.75 15.15 14.64
Ours (L) 10.39 11.45 11.66 18.09 18.42 15.35 22.26 22.52 21.75
Ours (Q) 14.32 12.09 14.95 19.29 15.42 15.95 22.77 23.12 20.46
Table 1: Fix Rate (FR%) on CelebA for varying numbers of added images, models, and methods. Ours (L) and Ours (Q) denote LLaVA-v1.5-7B and Qwen2.5-VL-7B as the large vision–language model filters, respectively. The highest FR in each column is marked in bold.
Method ResNet (base acc: 71.73%) ViT (base acc: 97.42%) CLIP (base acc: 93.78%)
DA-CDM 5.24 7.32 8.53 6.20 13.18 13.57 2.89 4.34 5.95
Mask-ControlNet 1.73 2.87 2.33 -6.59 10.47 3.88 -5.47 1.61 -1.29
HiBug_Class 1.80 1.49 0.74 -11.63 -13.57 -5.43 2.25 6.91 0.80
HiBug_Task 6.30 4.07 5.77 9.30 13.95 6.20 -7.88 6.43 5.95
Ours (L) 8.38 8.70 7.75 30.62 29.07 26.74 11.41 16.40 10.61
Ours (Q) 9.13 6.01 11.14 31.01 25.97 38.37 12.70 17.68 18.81
Table 2: Fix Rate (FR %) on ImageNet10 under varying image counts, models, and methods.
@@ 5.1 Experimental Setup
Datasets. We evaluate our method on two classification tasks that exhibit attribute-based failure modes and we use an 8:1:1 split for training, validation, and testing, respectively.
Lipstick-wearing classification. We use the CelebA dataset [liu2015faceattributes] and follow the same data split protocol as [chen2023hibug] (80,000, 10,000, 10,000 for train/val/test). The task is to predict whether a person is wearing lipstick, a label known to be correlated with other attributes such as gender and hair color. In the following experiments, we refer to this dataset simply as CelebA.
ImageNet-10 classification. We construct a 10-class subset of ImageNet [deng2009imagenet] containing the following categories: backpack, barber chair, coffee mug, desk, electric guitar, park bench, pitcher, purse, rocking chair, and water bottle. Each class contains 1,300 images. This subset is selected to study classification failures related to visual attributes such as texture and color. In the following experiments, we refer to this dataset simply as ImageNet10.
@@
  • DA-CDM  [fang2024data] is a data augmentation method for object detection. It uses a controllable diffusion model guided by visual priors from original images, which enables direct reuse of existing bounding box annotations. It then applies a category-calibrated CLIP score to filter generated data and ensure high-quality, text-aligned samples.
  • Mask-ControlNet  [huang2024data] is a text-guided image generation pipeline that uses ControlNet with facial occlusion masks to synthesize diverse face images under specific occlusions as a data augmentation method to improve model robustness.
  • HiBug_Class  [chen2023hibug]. A Class-level method that augments training data with synthetic images. For each class, it uses a diffusion model with a prompt:
@@
  • HiBug_Task [chen2023hibug]. A Task-level variant of HiBug_Class that targets failure-prone attributes. It selects attribute slices with the highest validation error and generates prompts to guide a diffusion model:
@@
    Note: Unless otherwise specified, “HiBug” refers to this optimized “HiBug_Task” variant in the paper. We do not compare with HiBug2 [chen2025hibug2], which is a data selection method and differs from the synthetic generation strategy.
Metrics. We focus on targeted improvements for rare slices, which are critical for fairness and safety, while maximizing overall classification accuracy. Following prior work [lai2023improvingpredictionbackwardcompatiblilitynlp], we report the Fix Rate (FR), defined as
$FR=\frac{Acc_{\text{after}}-Acc_{\text{before}}}{1-Acc_{\text{before}}},$ (6)
which measures the fraction of previously misclassified samples that are corrected. Here, $Acc_{\text{after}}$ denotes the accuracy obtained after applying the proposed method, and $Acc_{\text{before}}$ represents the baseline accuracy from standard training using the vision model.
Implementation Details. All experiments use an NVIDIA A100 GPU. We take the CelebA “wearing lipstick” classification task as an example. We evaluate three backbone architectures for this classification task: ResNet-18 [he2016deep], ViT-B/16 [dosovitskiy2020vit], and CLIP (ViT-B/32) [radford2021learning], all of which are initialized with random weights. All vision models are trained using cross-entropy loss. For rare-case bug discovery, we set the rarity threshold $\rho=0.05$ and the accuracy difference threshold $\epsilon=0.03$. For synthetic augmentation, we use ControlNet [zhang2023adding], a CDM based on Stable Diffusion 1.5 [rombach2022high] and conditioned on soft HED boundaries [xie2015holistically], with 30 DDIM inference steps. Soft HED boundaries preserve structural details, making this approach suitable for attribute-preserving edits like recoloring and stylizing. To filter generated images, we employ the LVLMs Qwen2.5-VL-7B [bai2025qwen25vl] and LLaVA-v1.5-7B [liu2023improvedllava]. Generating 1,000 images with ControlNet takes about one hour, and filtering these 1,000 images with the LVLM takes ten minutes. Filtering accuracy for most attributes (e.g., hair color, skin tone) exceeds 90%, which is consistent with the human audit in Section 5.6. Combining three attributes yields at least a 70% pass rate for filtered images, showing that the diffusion model produces high-quality samples and that the LVLM filtering is reliable.
@@ 5.2 Main Results
On CelebA, we select rare-case bugs defined by attribute–value combinations red hair, brown skin, and sad emotion for ResNet and ViT, and yellow hair, brown skin, and sad emotion for CLIP, based on the most frequent patterns identified among failure slices. SafeFix consistently achieves the highest test accuracy, i.e., the highest FR, across all models (ResNet, ViT, and CLIP) and varying levels of synthetic augmentation. For example, with 1,000 added images, our method improves FR by +14.32% (ResNet), +10.29% (ViT), and +22.77% (CLIP) relative to their base accuracies. These results show that our attribute-targeted augmentation and filtering pipeline is effective in repairing rare-case failure slices, outperforming both CDM-based and HiBug baselines.
@@
On ImageNet10, similar trends emerge, as shown in Table 2. For ResNet and ViT, we target rare-case bugs involving pink color and fabric texture, while for CLIP we use orange color and fabric texture. Across all models, our proposed method consistently surpasses the baseline methods. Specifically, our method improves ResNet accuracy by +11.14% (with 1,000 images) compared to the base model. ViT and CLIP also exhibit a steady improvement compared to other methods. Table 3 shows that the proposed SafeFix also achieves better overall accuracy on ImageNet10 compared with the baselines.
Method ResNet
100 Images 500 Images 1k Images
Base 71.73
DA-CDM 73.21 73.80 74.14
Mask-ControlNet 72.22 72.54 72.39
HiBug_Class 72.24 72.15 71.94
HiBug_Task 73.51 72.88 73.36
Ours (L) 74.10 74.19 73.92
Ours (Q) 74.31 73.43 74.88
Table 3: Test accuracy (%) on ImageNet10 using ResNet-18.
@@
Analysis. Combined with Tables 1, 2, 3, and all other test-accuracy results in Appendix E, our experiments show that baseline performance is often unstable. For instance, baselines like DA-CDM and Mask-ControlNet show some improvement on object and facial attribute tasks, respectively, due to their use of conditional diffusion models. However, they are not targeted model repair methods, so their overall performance is inferior to both HiBug and SafeFix. This lack of a targeted strategy means their gains are task-specific and not robustly transferable. Similarly, the accuracy of HiBug does not improve substantially. This is likely because it does not precisely target failure-critical attributes, instead applying general augmentations that can introduce noise or modify unintended features. As shown in Figure 2, this issue is compounded by the lack of LVLM filtering, resulting in generated data that is often misaligned with the intended rare-case fixes.
In contrast, our method (SafeFix) consistently shows stable or improving performance. This robustness indicates our attribute-targeted augmentation and LVLM filtering are highly effective at correcting failure-prone subpopulations with meaningful data.
@@ 5.3 SafeFix Can Effectively Fix Rare-case Bugs
Figure 3: Accuracy comparison of ResNet-18 trained using different augmentation methods. The dashed line represents the average overall accuracy without additional synthetic training data.
To verify that SafeFix’s improvements specifically address targeted rare-case bugs rather than merely enhancing overall performance, we analyze attribute-level validation accuracy changes on CelebA and ImageNet10, as shown in Figure 3. For clarity, the ImageNet10 plot includes only three representative attributes—color, background, and texture—as the dataset contains many attribute dimensions. Specifically, the left plot highlights improvements for CelebA after adding 5,000 synthetic images to the original 80,000 training samples. The right plot demonstrates accuracy gains on ImageNet10, achieved by adding 100 synthetic images to the original training set of 10,400 samples.
Taking ImageNet10 as an example, our targeted synthetic augmentation on pink color and fabric texture significantly improved accuracy for these selected rare-case attributes. Accuracy for the pink color attribute increased from 69.90% to 74.76%, surpassing the overall accuracy of 74.31%. Similarly, accuracy for the fabric texture attribute improved from 65.85% to 75.61%, also exceeding the overall accuracy. In contrast, attributes not explicitly targeted by augmentation, such as the rocks background, exhibited minimal or no improvement—its accuracy remained unchanged—highlighting that the gains from SafeFix are concentrated on the intended rare-case bugs rather than uniformly distributed across all attributes.
These substantial attribute-specific improvements confirm that SafeFix effectively repairs identified rare-case failure slices rather than providing a generalized performance boost. SafeFix also shows that all attributes improve across both datasets and introduces no new bugs, which indicates that the method remains stable outside the targeted regions. Attributes not targeted by augmentation show negligible accuracy changes, further reinforcing that SafeFix precisely and safely addresses the targeted rare-case failures.
@@ 5.4 SafeFix Directly Addresses Diagnosed Failure Modes, Not Merely Augments Data
To verify that SafeFix’s performance gains stem from accurately fixing rare-case bugs rather than from generic data augmentation, we compare red-hair and yellow-hair augmentations on CLIP. As discussed in Section 5.2, we select red hair for ResNet and yellow hair for CLIP in the CelebA dataset, based on which attribute is more likely to trigger rare-case bugs. For CLIP, both red hair and yellow hair are low-frequency attributes that meet the rarity criterion in Eq. (4). However, only the yellow hair slice additionally satisfies the low-accuracy criterion in Eq. (5), making it a true rare-case bug slice for CLIP.
Figure 4: Test accuracy(%) improvements for red-hair vs. yellow-hair augmentation on CLIP.
Figure 4 confirms that augmenting 1000 images using yellow-hair samples—the diagnosed failure mode for CLIP—substantially improves accuracy on its target slice (+4.51%). Furthermore, this targeted augmentation provides positive gains across all other attributes, including on the "Red Hair" slice (+0.91%) and a significant boost to "Overall" accuracy (+2.66%). Conversely, augmenting with red hair, which is not a diagnosed bug, reveals a harmful outcome. Most notably, it degrades performance on its own target slice by -0.79%. This counter-intuitive result contrasts with the positive gains seen for ResNet (Figure 3(a)), highlighting that different architectures can react to synthetic data in unpredictable ways. While red-hair augmentation does provide minor gains on other attributes (e.g., +0.75% on sad emotion), its minimal overall accuracy impact (+0.62%) and significant negative side effect on the targeted class itself underscore the importance of our diagnosis-driven approach. Augmenting the correctly diagnosed bug (yellow hair) leads to effective and safe repair, while augmenting an undiagnosed attribute (red hair) can be ineffective and harmful.